Two 25 kg children are sitting on a merry-go-round that is rotating at 0.10 rad/s. If one child is 1.0 m from the center and the other is 1.5 m from the center, what is their combined angular momentum
Answer #1

According to given data,

The angular momentum is,

L=moment of Inertia * angular velocity

L=I*w=(mr²)*w

So , the moment of Inertia of one child,

I1=m(r1)²

I1=25*(1)²=25 kg m²

And moment of Inertia of other child,

I2=m(r2)²

I2=25*(1.5)² =56.25 kg m²

Now, combined angular momentum,

L=total moment of Inertia *angular velocity

L=(I1+I2)*w

L=(25+56.25)*0.10=81.25*0.10

L=8.125 kg m²/s